Two concrete spans of a 170 m long bridge are
placed end to end so that no room is allowed
for expansion. If the temperature increases by 20◦C, what
is the height to which the spans rise when
they buckle? Assume the thermal coefficient
of expansion is 1.2 × 10−5
(◦C)−1.Answer in units of m.
1
Expert's answer
2020-03-10T11:18:43-0400
"L_i=0.5L=0.5(170)=85\ m"

After expansion, the length of one span is


"L_f=L_i(1 +αΔT)=85(1 +(1.2 \cdot10^{−5})(20))=\\=85.0204\ m"

From the Pythagorean theorem,


"h=\sqrt{L_f^2-L_i^2}"

"h=\sqrt{85.0204^2-85^2}=1.86\ m"